A shaft subjected to bending moment varying from -200 N m to +500 N m and a varying torque from 50 N m to 175 N m. If material of the shaft is 30C8, stress concentration factor is 1.85, notch sensitivity is 0.95 reliability 99.9% and factor of safety is 1

  1. A shaft subjected to bending moment varying from -200 N m to +500 N m and a varying torque from 50 N m to 175 N m. If material of the shaft is 30C8, stress concentration factor is 1.85, notch sensitivity is 0.95 reliability 99.9% and factor of safety is 1.5, find the diameter of the shaft. The load on a bolt consists of an axial pull of 10 kN together with a transverse shear force of 5 kN. Find the diameter of bolt required according to 

  1. Maximum principal stress theory;

  2. Maximum shear stress theory; 

  3. Maximum principal strain theory;

  4. Maximum strain energy theory; and

  5. Maximum distortion energy theory. 

Take permissible tensile stress at elastic limit = 100 MPa and poisson’s ratio = 0.3
Answer #1

obtain the property of 300g material from appendix of о aney text book. Tensile strength is (Sut) = 500 N/mm? Yield strengthTm Tmax + Tmin 2 50 +175 2 11205 Nm Calculate alternate topque using the relation Ta Tmax -Tmin २ 175-50 2 = 6205 nom ben Cal327 350 X1000 N-mm IT d 3 3, 565, 070-725 d3 Calculate Mean and alternate shear stresses Pm 16 Tom Td3 164 112-581000 N-mm Td(a), -/ (3,565,070 070.125) + 3 (318, 30909 •g) d? d3 3, 607, 449.63 N/mm² d3 (3) m² + 3 Tont e ro 1527887.45 92.06) * + 3 (a (gut) b For Machined a = 4.51 b=-0.265 ka = 4:51 (50050-265 8688 kca reliability factor at aqoq.4. Reliability ke=0.753 KbDetermine actual stress con centration factor (kg) using the relation q 인 kf-1 KL-1 Here, q is notch sensitivity, if actual s6586:13235 4-6909.82 ㅜ t d3 L ㅗ 1.5 d3 da = 43.1324 mm. Therefore, diameter of shaft is 43.13 mmGiven Data! Axial load bolt (P) TO KN Tranverse load 5kN on bot (vo Permisible tensile stress at Elastic limit (pes) - 100MPa러 = A Here, ? is Shear stress in bolt and Vis Shear force. substitute 5KN for v and Id2 for A. 퓨어 a 구 5 kN XON/IKN 포 6366,2 러1,2 (5) I ſE² + 92 12732.4 (2732.4 :) +(636602 dz 2d2 d2 2 6366.2 d2 2 9003.16 d2 ハ - 15369.4 N/mm² 02 a 2636.96 N/mm² DetermDetermine diameter of bolt according to Maximum shear stress theory using the relation a02 о per 15369.4 2636.96 + 100 N/mm dDetermine diameter of bolt according theory using strain energy the Gelation. 22+ 22 zu oara oper > 5369.4 +) + (-263696 + 2416843:41 - 100 d2 d 12.978 mm to Therefore, diameter of bolt according distortion energy theory is 112.978 mm